When Amy rollerskates, she moves 120 yards per minute. What is her speed in miles per hour?
Found 2 solutions by mananth, oberobic:
Answer by mananth(16946)

120 yards/minute
=120 yards/60 seconds
= 2 yards/second
1 yard = 1760 mile
1second = 1/3600 hours
2 yards/ second ==> (2/1760)/(1/3600)= (2/1760)*(3600/1)
=4.09 mph

Answer by oberobic(2304)

120 yd/min = x mi/hr
.
120 yd/min * 60 min/hr = 7200 yd/hr
.
5280 ft/mile / 3 ft/yd = 1760 yd/mi
.
7200 yd/hr / 1760 yd/mile = 4.0909 mile/hr